# Spectral Architecture Search for Neural Network Models

### v1

@@ 1 Introduction
Neural networks are very effective machine learning tools that prove extremely valuable in unwinding the best representation of the data at hand. To improve the ability of neural networks to automatically perform the tasks assigned, innovative architectures have been proposed and thoroughly tested. Employed architectures have been customarily developed by human experts, with manual, time-consuming, and error-prone processes. To go beyond manual design, novel algorithmic strategies for automated discovery of optimal neural architectures have been developed. Consequently, architecture engineering has become a relevant field of active research [1, 2].
@@
Neural Architecture Search (NAS), the process that seeks to optimize network architecture, has been successfully applied on tasks as image classification [3, 4], object detection [3], or semantic segmentation [5], yielding remarkable performance, as compared to manually designed benchmarks. According to [1], NAS is a subfield of Automated Machine Learning (AutoML) [6], the process that aims at automating the steps propaedeutic to applying machine learning to real-world problems. It also shows a notable overlap with hyperparameter optimization, a critical process in machine learning that involves selecting the optimal set of hyperparameters for a learning algorithm.[7] and meta-learning, a subfield of machine learning that focuses on training models to understand and adapt to new tasks autonomously [8].
@@
Starting from these premises, we here propose SPARCS (SPectral ARchiteCture Search), a training algorithm for deep neural networks which builds on the spectral parametrization of the inter-layer transfer matrices to automatically identify the best possible architecture given the specific problem under scrutiny. To this end, the multi-layered network is initially set to operate as a perceptron: only direct connections from the input nodes towards the destination layer are hence active. Upon training, and depending on the inherent complexity of the examined problem, hidden layers get recruited, via feedforward or skip (long ranged) connections, and thus contribute to the performed computation. As we shall prove, this yields a rather compact network, in terms of links activated ex post, as compared to viable approaches rooted on a conventional (non spectral) formulation of the training problem. Importantly, SPARCS explores the space of possible architectures by scanning a continuous and differentiable manifold, with gradient-based optimization algorithms. In this work, we set up the theory of SPARCS for a neural network of arbitrary dimensions and validate its adequacy against two simple benchmark applications. The first aims at testing the efficacy of our model in grasping the linear to non-linear transition of a dataset, by adequately adapting the architecture of the learned model. The second application deals with a regression problem of a randomly generated neural network with weight distributions which can be approximated with just one hidden layer. This solution is indeed correctly spotted by our method.
@@
Historically, NAS methods leveraged evolutionary algorithms to explore the architecture’s space. One famous example is NEAT (Neuro-Evolution of Augmenting Topologies) [9]. This method encodes the topology of the network in a set of genes, then a population of networks is created and put under evolutionary pressure. The fitness of a network is given by the displayed performance on the examined task (classification, regression, etc.). This can be seen as a sort of hyperparameter optimization process, where the search space of parameters is explored through evolutionary means. Another family of methods reformulate NAS as a veritable hyperparameter search problem. These are the so called RL-NAS (Reinforcement Learning based Neural Architecture Search). These latter methods use a neural network to generate a neural network architecture, then train and test an instance of such architecture on the task at hand, and finally use the recorded performance as a reward signal to feed the generating network scheme[10, 11]. These two families of methods are still used and researched today, however, with time, performance considerations pushed the community towards adopting a differentiable architecture search approach: during training, the optimisaztion of both the architecture and the model’s weights are simultaneously carried out. One famous example is DARTS (Differentiable ARchiTectures Search). The original algorithm spawned then popular variants such as PC-DARTS, GDAS, FairDARTS [12, 13, 14, 15]. The main idea behind all of these methods is to associate continuous parameters to the topological features of the network, and then optimize them, along with the network’s weights, through iterative procedures, like gradient descent. In this context SPARCS places itself as a new differentiable NAS method, but with three notable addition that we anticipate hereafter:
@@
More into details, SPARCS exploits a non trivial generalization of the so called spectral parametrization [16, 17, 18, 19], a recently introduced scheme for handling the optimization of feedforward networks. In a nutshell, feedforward neural networks can be pictured as an ordered sequence of linear transformations (interspersed by non linear filters), each encoded in a square adjacency matrices for the ensuing bipartire directed graph. The spectral paradigm reformulates the optimization process in dual space, the eigenvalues and the eigenvectors of the inter-layer adjacency matrix, supplemented with the inclusion of self-loops, acting as target free parameters. To be able to operate Neural Architecture Search in the spectral domain, one needs first to extend the realm of validity of the spectral parametrization beyond the simple entry setting, where the information transfer between two consecutive layers is solely considered. Working along these lines, we will here set the mathematical foundation for a spectral characterization for the multi-layered interactions between an arbitrary collection of inter-tangled stacks, of different sizes and subject to local and long-ranged couplings. This generalized framework prompts to a wider usage of the spectral technology, beyond what so far explored in the literature and aside the specific application here discussed.
@@ 2 Mathematical background: the spectral theory for a feed-forward network with local connections
Consider a deep feedforward network made of $\ell$ distinct layers and label each layer with the index $i$ $(=1,...,\ell)$. Denote by $N_{i}$ the number of neurons which belong to the $i-$th layer. The total number of parameters that one seeks to optimise when dealing with a typical fullyconnected neural network (all neurons of any given layer with $i<\ell-1$ are linked to every neurons of the adjacent layer) equals $\sum_{i=1}^{\ell-1}N_{i}N_{i+1}$, when omitting additional bias. At this level, no skip connections, namely long-ranged links between distant (i.e. non adjacent) layers are considered. Working within this setting, we will review in the following the core idea of the spectral parametrization as introduced in [16]. This prelude is meant to provide the necessary background for the non trivial extension hereafter addressed.
@@
$k$$k+1$$k+2$
$k$$k+1$$k$$k+1$$k+1$$k+2$$k+1$$k+2$
$=$$A^{(k)}$$\textbf{a}_{k}$$\textbf{a}^{\star}_{k}$$=$$A^{(k+1)}$$\textbf{a}_{k+1}$$\textbf{a}^{\star}_{k+1}$
Figure 1: The linear information transfer across adjacent layers is encoded in a squared adjacency matrix. Notice that the transformation from $\textbf{a}_{k}^{\star}$ to $\textbf{a}_{k+1}$ is a projection, mathematically implemented via a suitable block matrix $\pi$, with all the entries set to 00 except for those referred to the upper right $N_{2}\times N_{2}$ square block, which are set to the identity.
@@
$k$$k+1$$k$$k+1$$k$$k+1$$k$$k+1$
$=$$A^{(k)}$$\textbf{a}_{k}$$\textbf{a}^{\star}_{k}$$=$$A^{(k)}$$\textbf{a}_{k}$$\textbf{a}^{\star}_{k}$
@@
(b) Eigenvalue matrix $\Lambda$. The wavy diagonal line indicates elements that generally have values different from 00 or $1$.
[…]
(c) Adjacency matrix $A$ produced by the spectral method.
@@ 3 A simple network with just one hidden layer
[…]
@@ 4 Extending the spectral theory to deep networks of arbitrary size
The effect of such activation is presented in panel (b) of Figure 6, for an hypothetical setting of exclusive pedagogical value. A new bundle of connections is activated between layer $1$ and $B+1$ via layer $j$, which impacts onto the generated output, provided the networks is operated with non linear modality (multiple applications of $A_{B}$, interposed by the application of suitable non linear filters). Another interesting effect is the implicit regularization that occurs by initializing the eigenvalues to zero. This latter induces a gradual increase in rank and progressively allows to explore more complex solutions [20]. Remarkably, as we will demonstrate in the experiments section, this approach also prioritizes solutions with smaller depths without requiring explicit regularization targeted to that aspect. The spectral parametrization materializes therefore in a natural strategy to modulate the functional complexity of the computing network (expressed in terms of non linear elaboration) that is differentiable and trainable. Motivated by the above, we will introduce an apt regularization factor designed so as to enhance the non linear elaboration power of the computing network, only when mandatorily needed. Relevant non linear transformations will be self-consistently enforced during training stages, depending on the inherent intricacies of the task being explored. This observation paves the way to a veritable Architecture Search strategy as we shall outline in the forthcoming Section with reference to a simple test case model.
@@ 5 Spectral Architecture Search: simple demonstrative examples
Figure 8: Number of parameters needed (assuming constant layer dimension equal to $100$) to perform architecture search against the the total number of employable layers (not just the number of hidden layers). The scaling indicates that the spectral formalism (orange line) requires significantly fewer parameters, except for the two-layer case, where direct parameterization (blue curve) has a slight advantage.
@@ 6 Conclusions
By treating a multilayer perceptron as a multipartite graph, we established an analytical correspondence between the links of any two layers (including not only adjacent interlayer connections but also the so-called ”skip-layer” connections) and the eigenvalues and eigenvectors of the graph’s adjacency matrix. This parameter-sharing mechanism is induced by such an analytical correspondence as formalized in Eq. (4.3).
@@ Appendix B Explicit expressions for the blocks of $\Phi_{B}^{-1}$
For $n+1<i-j$, we obtain $\mathbb{O}$, as desired. Conversely, if $n+1\geq i-j$ yields:
@@ Appendix C On the block structure of $A_{B}$
and this ends the proof. Notice in fact that for $i-j=1$ the last factor of this latter expression is not present. Equation (4.3) is thus formally recovered.
@@ Appendix D Proof of binomial formulae
$-\sum_{k^{\prime}=1}^{B}(-1)^{k^{\prime}}\binom{B}{k^{\prime}}=-\sum_{k^{\prime}=1}^{B}(-1)^{k^{\prime}}\binom{B}{k^{\prime}}-\binom{B}{0}+\binom{B}{0}$ (D.9)
@@
Appendix E Experiments
@@
Figure 11: Dataset for $\beta=5$ and $\alpha$ varying in the range from 00 to $1$.
@@
We trained on two sets of datasets, one with $\beta=5$ and one with $\beta=10^{3}$, for a total of $42$ datasets. For each dataset we trained $100$ model samples. An example of the goodness of fit for $\beta=5$ and model sample number 00 is shown in figure 12.
Figure 12: Fit data for model 00 on the datasets with $\beta=5$. In red we have shown the model’s outputs for different inputs, and in green we report the error associated to the model with respect to the true target function $f$.
@@
Given this definition if the norm of $\Gamma$ is equal to 00, no signal is being processed by the hidden layer, and all the information is handled by the linear skip connection layer.

### v2

@@ 1 Introduction
Neural networks are very effective machine learning tools that prove extremely valuable in unwinding the best representation of the data at hand. To improve the ability of neural networks to automatically perform the assigned tasks, innovative architectures have been proposed and thoroughly tested. Employed architectures have been customarily developed by human experts, with manual, time-consuming, and error-prone processes. To go beyond manual design, novel algorithmic strategies for automated discovery of optimal neural architectures have been developed. Consequently, architecture engineering has become a relevant field of active research [1, 2].
@@
Neural Architecture Search (NAS), the process that seeks to optimize network architecture, has been successfully applied on tasks as image classification [3, 4], object detection [3], or semantic segmentation [5], yielding remarkable performance, as compared to manually designed benchmarks. According to [1], NAS is a subfield of Automated Machine Learning (AutoML) [6], the process that aims at automating the steps propaedeutic to applying machine learning to real-world problems. It also shows a notable overlap with hyperparameter optimization (a critical process in machine learning that involves selecting the optimal set of hyperparameters for a learning algorithm [7]) and meta-learning (a subfield of machine learning that focuses on training models to understand and adapt to new tasks autonomously [8]).
@@
Starting from these premises, we here propose SPARCS (SPectral ARchiteCture Search), a training algorithm for deep neural networks which builds on the spectral parametrization of the inter-layer transfer matrices to automatically identify the best possible architecture given the specific problem under scrutiny. To this end, the multi-layered network is initially set to operate as a perceptron: only direct connections from the input nodes towards the destination layer are hence active. Upon training, and depending on the inherent complexity of the examined problem, hidden layers get recruited, via feedforward or skip (long ranged) connections, and thus contribute to the performed computation. As we shall prove, this yields a rather compact network, in terms of links activated ex post, as compared to viable approaches rooted on a conventional (non spectral) formulation of the training problem. Importantly, SPARCS explores the space of possible architectures by scanning a continuous and differentiable manifold, with gradient-based optimization algorithms. In this work, we set up the theory of SPARCS for a neural network of arbitrary dimensions and validate its adequacy against simple benchmark applications. The first benchmark aims at testing the efficacy of our model in grasping the linear to non-linear transition of a dataset, by adequately adapting the architecture. The second deals with a regression problem of a randomly generated neural network, with weight distributions which can be approximated with just one hidden layer. This solution is indeed correctly spotted by our method. Additional tests performed by operating SPARCS in conjunction with Convolutional Neural Networks for image classification tasks (CIFAR-10 and CIFAR-100) are reported in Appendix F and commented upon in the main body of the paper. The hyperparameters employed in carrying out the above tests are summarized for convenience in a set of Tables reported in Appendix G.
@@
Historically, NAS methods leveraged evolutionary algorithms to explore the architecture’s space. One famous example is NEAT (Neuro-Evolution of Augmenting Topologies) [9]. This method encodes the topology of the network in a set of genes, then a population of networks is created and put under evolutionary pressure. The fitness of a network is given by the displayed performance on the examined task (classification, regression, etc.). This can be seen as a sort of hyperparameter optimization process, where the search space of parameters is explored through evolutionary means. Another family of methods reformulate NAS as a veritable hyperparameter search problem. These are the so called RL-NAS (Reinforcement Learning based Neural Architecture Search). These latter methods use a neural network to generate a neural network architecture, then train and test an instance of such architecture on the task at hand, and finally use the recorded performance as a reward signal to feed the generating network scheme[10, 11]. These two families of methods are still used and researched today, however, with time, performance considerations pushed the community towards adopting a differentiable architecture search approach: during training, the optimization of both the architecture and the model’s weights are simultaneously carried out. One famous example is DARTS (Differentiable ARchiTectures Search). The original algorithm resulted in popular variants such as PC-DARTS, GDAS, FairDARTS [12, 13, 14, 15]. The main idea behind all of these methods is to associate continuous parameters to the topological features of the network, and then optimize them, along with the network’s weights, through iterative procedures, like gradient descent. In this context SPARCS places itself as a new differentiable NAS method, but with three notable addition that we anticipate hereafter:
@@
As an additional remark, we stress that the above referenced NAS methods prove useful when the training for the specific problem under scrutiny requires a limited amount of time, in such a way that multiple training attempts can be tolerated at the price of a reasonable computation cost. Unfortunately that is not always the case. If the training takes several weeks on powerful computing resources, the usage of methods like NEAT and DARTS, which come along multiple training sessions, appear prohibitive, at least on standard hardware. SPARCS can, at least in principle, fill this gap as (i) the architecture search is performed during the main (and sole) training session, with no need for multiple runs; (ii) the number of trainable parameters is set by the largest architecture in the explored example. More into details, SPARCS exploits a non trivial generalization of the so called spectral parametrization [16, 17, 18, 19], a recently introduced scheme for handling the optimization of feedforward networks. In a nutshell, feedforward neural networks can be pictured as an ordered sequence of linear transformations (interspersed by non linear filters), each encoded in a square adjacency matrices for the ensuing bipartite directed graph. The spectral paradigm reformulates the optimization process in dual space, the eigenvalues and the eigenvectors of the inter-layer adjacency matrix, supplemented with the inclusion of self-loops, acting as target free parameters. Notice that the spectral approach to training requires no costly spectral decomposition step, as the network is formulated from the outset in terms of eigenvectors and eigenvalues. To be able to operate Neural Architecture Search in the spectral domain, one needs first to extend the realm of validity of the spectral parametrization beyond the simple entry setting, where the information transfer between two consecutive layers is solely considered. Working along these lines, we will here set the mathematical foundation for a spectral characterization for the multi-layered interactions between an arbitrary collection of inter-tangled stacks, of different sizes and subject to local and long-ranged couplings. This generalized framework prompts to a wider usage of the spectral technology, beyond what so far explored in the literature and aside the specific application here discussed.
Recapitulating one more time, we focus first on extending the validity of the spectral parametrization beyond the simple case of two consecutive layers, so as to account for the general setting where an arbitrary number of layers is accounted for. As an immediate application of the proposed framework, we present and thoroughly test an innovative scheme for automated neural network architecture search. The algorithm initializes the networks as simple perceptrons. Then, it adds additional complexity via hidden layers based on task demands, with a parsimonious approach which tends to favor the simplest solution possible, given the problems constraints. This results in efficient computational models with fewer parameters, as compared to other viable alternatives. Our conclusions are supported by a campaign of tests that we have performed on benchmark models. To contribute with a novel formulation of the optimization problem (which generalizes the existing spectral perspective) and to provide a novel scheme for architecture search (which exploits the proposed mathematical framework) define the main motivations of our study.
@@ 2 Mathematical background: the spectral theory for a feed-forward network with local connections
[…]
@@
kkk+1k+1k+2k+2
kkk+1k+1kkk+1k+1k+1k+1k+2k+2k+1k+1k+2k+2
$A^{(k)}$$\textbf{a}_{k}$$\textbf{a}^{\star}_{k}$$A^{(k+1)}$$\textbf{a}_{k+1}$$\textbf{a}^{\star}_{k+1}$
Figure 1: The linear information transfer across adjacent layers is encoded in a squared adjacency matrix. Notice that the transformation from $\textbf{a}_{k}^{\star}$ to $\textbf{a}_{k+1}$ is a projection, mathematically implemented via a suitable block matrix $\pi$, with all the entries set to 0 except for those referred to the upper right $N_{2}\times N_{2}$ square block, which are set to the identity.
@@
$kkk+1k+1kkk+1k+1$
$A^{(k)}$$\textbf{a}_{k}$$\textbf{a}^{\star}_{k}$$A^{(k)}$$\textbf{a}_{k}$$\textbf{a}^{\star}_{k}$
@@
$\mathcal{L}_{1}$$\mathcal{L}_{2}$$\Lambda=$$N_{1}$$N_{2}$$N_{1}$$N_{2}$
(b) Eigenvalue matrix $\Lambda$. The wavy diagonal line indicates elements that generally have values different from 0 or 11.
@@
(c) Adjacency matrix AA produced by the spectral method.
@@ 3 A simple network with just one hidden layer
$\mathcal{L}_{1}^{(2)}$$\mathcal{L}_{2}^{(2)}$$\mathcal{L}_{3}^{(2)}$$N_{1}$$N_{2}$$N_{3}$$N_{1}$$N_{2}$$N_{3}$$\phi^{(2)}_{1}\mathcal{L}^{(2)}_{1}-\mathcal{L}^{(2)}_{2}\phi_{1}^{(2)}$$\phi^{(2)}_{2}\mathcal{L}^{(2)}_{2}-\mathcal{L}^{(2)}_{3}\phi_{2}^{(2)}$$\Big{(}\mathcal{L}^{(2)}_{3}\phi^{(2)}_{2}-\phi^{(2)}_{2}\mathcal{L}^{(2)}_{2}\Big{)}\phi^{(2)}_{1}$$A_{2}=$
@@ 4 Extending the spectral theory to deep networks of arbitrary size
The effect of such activation is presented in panel (b) of Figure 6, for an hypothetical setting of exclusive pedagogical value. A new bundle of connections is activated between layer $1$ and $B+1$ via layer $j$, which impacts onto the generated output, provided the networks is operated with non linear modality (multiple applications of $A_{B}$, interposed by the application of suitable non linear filters). Another interesting effect is the implicit regularization that occurs by initializing the eigenvalues to zero. This latter induces a gradual increase in rank and progressively allows to explore more complex solutions [20]. Remarkably, as we will demonstrate in the experiments section, this approach also prioritizes solutions with smaller depths without requiring explicit regularization targeted to that aspect. The spectral parametrization materializes therefore in a natural strategy to modulate the functional complexity of the computing network (expressed in terms of non linear elaboration) that is differentiable and trainable. Motivated by the above, we can introduce an apt regularization factor designed so as to enhance the non linear elaboration power of the computing network, only when mandatorily needed. Relevant non linear transformations will be self-consistently enforced during training stages, depending on the inherent intricacies of the task being explored.
Before proceeding we elaborate further on this point. SPARCS enables the model to be initialized with the minimal possible network topology: this is accomplished by initially setting all eigenvalues to zero, except for those associated to the neurons that belongs to the output layer. During training, the optimization of the spectral parameters modulates existing weights while, at the same time, allowing for a self-consistent growth of the network structure. To favor convergence towards compact architectures, so as to minimize the associated computational cost during deployment, one can introduce a regularization factor which acts on the eigenvalues. More specifically, the eigenvalues associated with the hidden layers are subjected to a (L1 or L2) regularization. The eigenvalues referred to the output layer are instead left untouched, while those pertaining to the input layer are set to zero. In formulae we will consistently adopt the following training loss:
[…]
where $\bm{\lambda},\bm{\phi}$ respectively stands for the collection of eigenvalues and eigenvectors, and $\mathcal{D}$ identifies the elements of the training set. The loss is hence given by the sum of a standard loss (e.g. MSE loss, or categorical cross-entropy loss, depending on the task at hand) and a regularization term $\Omega$ that depends exclusively on the eigenvalues associated to the hidden neurons $\bm{\tilde{\lambda}}$. Here, $\rho$ is a hyperparameter that sets the strength of the imposed regularization. As we shall discuss in the following, this simple strategy reliably yields minimal network topologies across a wide range of learning tasks. This observation paves the way to a veritable Architecture Search strategy as we shall outline in the forthcoming Section with reference to simple test case models. Further experiments will be discussed in Appendix F.
@@ 5 Spectral Architecture Search: simple demonstrative examples
Figure 8: Number of parameters needed (assuming constant layer dimension equal to $100$) to perform architecture search against the total number of employable layers (not just the number of hidden layers). The scaling indicates that the spectral formalism (orange line) requires significantly fewer parameters, except for the two-layer case, where direct parameterization (blue curve) has a slight advantage.
@@
Two additional tests for the newly proposed algorithm are discussed in Appendix F The first set of experiments are run on CIFAR-10. The architecture is composed by a minimal CNN module, that we consistently train across the experiments. The output of the CNN is then fed to a feedforward network, to perform classification. This latter feedforward network is initialized as a perceptron (a linear classifier). As in the spirit of the other reported tests, the architecture can be modulated by SPARCS and the performance gets self-consistently boosted (as compared to a benchmark linear model run in parallel) by taking full advantage of the non-linear character of the feedforward classifier. Indeed, the simple CNN module that we have employed is unable to reduce the analyzed task to a linear classification problem and SPARCS finds it advantageous to turn the classification from linear to non linear. In the second experiment, we considered CIFAR-100, a rather more complex dataset, but now we employed a pre-trained state-of-the-art convolutional neural network architecture (EfficientNetV2-S) as a pre-filter. In this case the powerful convolutional module turns the data into a linearly separable problem and SPARCS keeps across fine-tuning the minimal linear architecture, as it follows by the imposed initialization, without recruiting additional layers of the feedforward classifiers. With reference to these applications, we report also an estimate of the involved computational costs and related memory usage. For further technical information, please refer to Appendix F.
@@ 6 Conclusions
By treating a multilayer perceptron as a multipartite graph, we established an analytical correspondence between the links of any two layers (including not only adjacent inter-layer connections but also the so-called ”skip-layer” connections) and the eigenvalues and eigenvectors of the graph’s adjacency matrix. This parameter-sharing mechanism is induced by such an analytical correspondence as formalized in Eq. (4.3).
@@ Appendix B Explicit expressions for the blocks of $\Phi_{B}^{-1}$
For n+1<ijn+1<i-j, we obtain $\mathbb{O}$, as desired. Conversely, if $n+1\geq i-j$ yields:
@@ Appendix C On the block structure of $A_{B}$
and this ends the proof. Notice in fact that for ij=1i-j=1 the last factor of this latter expression is not present. Equation (4.3) is thus formally recovered.
@@ Appendix D Proof of binomial formulae
$\sum_{k^{\prime}=1}^{B}(-1)^{k^{\prime}}\binom{B}{k^{\prime}}=-\sum_{k^{\prime}=1}^{B}(-1)^{k^{\prime}}\binom{B}{k^{\prime}}-\binom{B}{0}+\binom{B}{0}$ (D.9)
@@
Appendix E More details regarding the first experiment
@@
Figure 11: Dataset for $\beta=5$ and $\alpha$ varying in the range from 0 to $1$.
@@
We trained on two sets of datasets, one with $\beta=5$ and one with $\beta=10^{3}$, for a total of $42$ datasets. For each dataset we trained $100$ model samples. An example of the goodness of fit for $\beta=5$ and model sample number 0 is shown in figure 12.
Figure 12: Fit data for model 0 on the datasets with $\beta=5$. In red we have shown the model’s outputs for different inputs, and in green we report the error associated to the model with respect to the true target function $f$.
@@
Given this definition if the norm of $\Gamma$ is equal to 0, no signal is being processed by the hidden layer, and all the information is handled by the linear skip connection layer.
@@
Appendix F Additional tests: operating SPARCS in conjunction with Convolutional Neural Networks for image classification tasks
To further testing SPARCS, we conducted a series of experiments on benchmark image classification tasks, specifically the CIFAR-10 and CIFAR-100 datasets. The results are reported below. To anticipate the main conclusion, we will prove that SPARCS can effectively operate in conjunction with various complex architectures, trained from scratch or with a fine-tuning of pre-trained models. We would like to point out that all the measurements of the allocated VRAM are reported considering the training phase. During inference the model can be remapped in the direct space, using (4.2) and the amount of memory needed is substantially reduced.
F.1 Testing on CIFAR-10
The CIFAR-10 dataset consists of 60,000 32x32 color images organized in 10 distinct classes, with 6,000 images per class. The dataset is divided into 50,000 training images and 10,000 test images. To assess the performance of SPARCS in this context we implemented two different models, both equipped with the same minimal convolutional feature extractor block, but with different classification heads. Specifically the feature extractor for both models is composed of convolutional layers, batch normalization, ReLU activation, and max pooling. The first model uses a linear classifier, and is trained in direct space (CNN), while the second model employs a classifier that is initiated as linear, but can grow along training as modulated by SPARCS to become a full nonlinear multi-layer perceptron (SPARCS+CNN). The training is performed from scratch using the Adam optimizer with a learning rate of 0.001, and a batch size of 128. The training is conducted for 300 epochs. The model is tested on the validation set after each epoch. For SPARCS we deploy an eigenvalue regularization constant of $10^{-4}$, with exclusion of (i) the eigenvalues corresponding to the last (output) layer, which bear no influence on the topology of the network; (ii) the eigenvalues associated to the first layer, which are fixed to zero to ensure a one-to-one correspondence between neurons and eigenvalues. The training is performed on a single NVIDIA RTX A5500 GPU, and the memory allocated during training, as well as the time elapsed, are kept track of. The results, shown in Figure 16, indicate that the SPARCS model achieves a higher accuracy as compared to the linear counterpart. This proves again the effectiveness of SPARCS in dynamically evolving the network topology during training so as to improve the recorded performance. This augmented performance is achieved with a modest increase in resource consumption, as shown in table 1. The eigenvalues corresponding to the neurons in the hidden layer, initially set to zero, grow to progressively assume values of the same order of magnitude as that displayed by the eigenvalues associated to the last layer of the collection. In words, SPARCS is capable of adjusting its topology to better tackle the supplied classification task. Interestingly, we observe the formation of two clusters in the eigenvalues distribution across the hidden layer: a limited fraction of the neurons are thus selected for usage, so yielding a compact architecture that will have reduced memory footprint and low computational cost in deployment and exploitation. The eigenvalue distribution is shown in Figure 17.
Figure 16: Performance comparison of SPARCS and linear model on CIFAR-10 dataset. The first curve (blue) represents the accuracy of the standard CNN with a linear classifier, while the second curve (green) represents the accuracy of the SPARCS+CNN. The third curve (red) shows the performances of the pruned SPARCS+CNN: this is obtained by removing all the neurons in the hidden layer. Notice that the three above curves stay close in the initial phases of the training. This seems to imply that the training is initially focused on optimizing the weights in the linear classifier space. However as the training progresses, the SPARCS model begins to potentiate the topology of the network, leading to a significant increase in the recorded performance.
Model GPU [^1] Memory allocated during training. (MB) Training Time (s) Test Accuracy (%)
CNN 17.9 1357.0 68.2
SPARCS+CNN 21.2 1361.98 73.9
Table 1: Comparison of resource consumption and performance resulting from the training of the SPARCS+CNN and the standard CNN on CIFAR-10 dataset. Notice how the SPARCS model achieves a significantly higher accuracy with only a modest increase in GPU memory consumption and training time.
(a) Eigenvalue distribution of the hidden layer in the SPARCS+CNN.
(b) Eigenvalue distribution of the last layer in the SPARCS+CNN.
Figure 17: Eigenvalue distributions of the SPARCS+CNN on CIFAR-10 dataset. The left figure shows the eigenvalue distribution of the hidden layer, while the right figure shows the eigenvalue distribution of the last layer. Notice that just a limited fraction of the eigenvalues associated to the hidden layer differ significantly from zero. This amounts to say that the algorithm has spontaneously selected a somehow minimal (or compact) architecture, by recruiting just a finite subset of active neurons in the hidden layer.
F.2 Testing on CIFAR-100
The CIFAR-100 has a similar structure to CIFAR-10, but it contains 100 classes with 600 images per class. In what follows, the same experimental setup as for CIFAR-10 is used, except for the adoption of a powerful pre-trained CNN architecture. Although the architecture used after the backbone is comparable to the previous one, this test shows the usability of SPARCS in realistic contexts and tasks, and its ability to operate with the data distribution downstream of a feature extraction. Tests integrating the method inside the full-architecture, with consequent ‘from scratch’ training, are left or upcoming work. The employed backbone is EfficientNetV2-S, a competitive convolutional neural network architecture. For both examined models, we fine-tune the pre-trained EfficientNetV2-S model, upon replacing the embedded output classifier with (i) a new linear module for the reference benchmark model and (ii) a SPARCS classifier for the competitor scheme. The fine tuning (one pass through the CIFAR-100 training set) is performed with the same hyperparameters employed for CIFAR-10, and in particular with the same regularization constant set to $10^{-4}$. Given the heavy amount of preprocessing going on in EfficientNetV2-S (and also since the original state-of-the-art like EfficientNetV2-S model is already equipped with a linear classifier), one can guess that the potential benefits as stemming by the activation of just one additional non linear layer in the classifier SPARCS section, would be negligible. Remarkably SPARCS, operated only on the train set, automatically finds that there is no need to augment the topology of the network to cope with the assigned task, given the massive pre-processing of the data. Hence, the linear initialization is kept unchanged through SPARCS. In table 2 we show the resource consumption and performance of the SPARCS EfficientNetV2-S model as compared to the standard EfficientNetV2-S model.
Model GPU [^2] Memory allocated during training. (MB) Fine Tuning Time (s) Test Accuracy (%)
EffNet 408.4 138.8 73.2
SPARCS EffNet 518.5 139.5 72.7
Table 2: Comparison of resource consumption and performance resulting from the fine-tuning of EfficientNetV2-S (EffNet) to CIFAR-100. Notice how both fine tuning paradigms achieve a similar performance: no advantage is found in recruiting more neurons from the hidden layer in the SPARCS classifier module.
(a) Eigenvalues of the hidden layer in the SPARCS EfficientNetV2-S.
(b) Eigenvalue distribution of the last layer in the SPARCS EfficientNetV2-S.
Figure 18: Eigenvalue distributions of the SPARCS EfficientNetV2-S on CIFAR-100 dataset. The left figure shows the eigenvalue distribution of the hidden layer, while the right figure shows the eigenvalue distribution of the last layer. Notice how the eigenvalues of the hidden layer are all orders of magnitude smaller than the eigenvalues of the last layer. This implies that SPARCS model did not call for an augmented topology, as it could have been anticipated given that the EfficientNetV2-S is already complex enough for handling the CIFAR-100 dataset.
Appendix G Hyperparameters used
For what concerns the regularization, in the experiments carried out, we have used both L1 and L2 norms as weight penalties. In order to further asses the flexibility of the method, we also varied the regularization strength parameter, and SPARCS appears to work well for a reasonable wide range of values (ranging approximately from $10^{-3}$ to $10^{-5}$). For many tasks however the sweet spot seem to be located around $10^{-4}$, following heuristics that its value should be around $10^{-3}$ of the Data Loss. In the following we present the parameters used for each task in the associated table.
Hyperparameters for the Simple Regression Experiment
Hyperparameter Value
Learning Rate $10^{-3}$
Batch Size 100
Number of Epochs 300
Max Number of Hidden Neurons 300
Optimizer Adam
Regularization Type L2
Regularization Strength $10^{-4}$
Activation Function ReLU
Loss Function MSE
NumPy Seed 42
Number of Trials 100
Table 3: Hyperparameters used in the first experiment. The numpy seed is used in the data generation phase, while number of trials refers to the number of repetition needed to give an estimate of the error bars in the results (standard deviations).
Hyperparameters for the Teacher-Student Experiment.
Hyperparameter Value
Learning Rate $10^{-3}$
Batch Size 1024
Number of Epochs 180
Number of Neurons in Teacher Hidden Layer 20
Max Number of Neurons in Student Hidden Layers 200
Optimizer Adam
Regularization Type L2
Regularization Strength $3\cdot 10^{-3}$
Activation Function ReLU
Loss Function MSE
NumPy Seed 42
Table 4: Hyperparameters used in the second experiment. The numpy seed is used in the data generation phase.
Hyperparameters for the CIFAR 10 and CIFAR 100 Experiments
Hyperparameter Value
Learning Rate $10^{-3}$
Batch Size 128
Number of Epochs 300
Max Number of Hidden Neurons 526
Optimizer Adam
Regularization Type L1
Regularization Strength $10^{-4}$
Activation Function ReLU
Loss Function Log Loss
Number of Trials 100
Table 5: Hyperparameters used in the third and fourth experiments (in appendix). Notice that all the parameters have been kept equal between the two experiments, except for the number of epochs: since on CIFAR 100 we only fine tuned, instead of training from scratch, we performed a single pass trough the training data. Also notice that with the term log loss we of course refer to the standard Categorical Cross-Entropy. Once again the number of trials line refers to the error estimation via empirical standard deviation measurements.